Two parallel plate air capacitors each of capacitance C were connected in series to a battery with emf E. Then one of the capacitors was filled up with uniform dielectric with relative permittivity k. What amount of charge flows through the battery?

Grade:11

2 Answers

Khimraj
5 years ago
initial charge Qi = 2CE
final  capacitance are C and KC
equivalent capacitance = KC/(1+K)
final charge = KCE/(1+K)
charge flown through the battery = final charge – initial charge
= KCE/(1+K) – 2CE
Gautam Jain
5 years ago
 
initial equivalent charge Qi = CE/2 
final  capacitance are C and KC
equivalent capacitance = KC/(1+K)
final charge = KCE/(1+K)
charge flown through the battery = final charge – initial charge
= KCE/(1+K) – CE/2
=1/2CE(K-1/K+1)